At a walk-in clinic, 120 patients arrive per hour on average. What is the probability that the receptionist needs to wait at least 2 minutes for the next patient to walk in? Round your answer to 4 decimal places.
Answer #1

expected number of patients in 2 minutes =120*2/60=4

therefore P( wait at least 2 minutes for the next patient to walk in) =P(no patients) =P(X=0)=e-4

=0.0183